# Do MLLMs Really Understand the Charts?

### v1

@@ Introduction
Methods Synthetic Charts Real Charts Overall
Box Area Radar Scatter Bar Line Combo Bar Line Combo
Open-source Models
InternVL3-2B (Zhu et al. 2025) 25.27 8.54 9.62 25.16 51.68 43.05 34.55 43.75 34.78 32.00 32.98
Qwen2.5-vl-3B (Bai et al. 2025) 46.15 14.63 17.95 52.30 72.35 80.79 62.65 66.96 53.04 46.40 56.62
Ovis1.6-llama3.2-3B (Lu et al. 2024) 12.09 3.66 7.69 14.66 13.13 10.93 11.69 8.04 16.52 12.00 11.66
Gemma-3-4B (Team et al. 2025) 9.89 3.05 12.82 11.82 9.22 12.25 6.81 8.04 12.17 7.20 9.34
Qwen2.5-vl-7B (Bai et al. 2025) 69.23 20.12 19.87 61.93 74.02 83.77 67.71 49.11 53.04 46.40 60.74
InternVL3-8B (Zhu et al. 2025) 36.73 12.80 12.18 39.17 43.58 47.68 36.82 38.39 46.09 34.40 36.73
Close-source Models
GPT-4o (OpenAI et al. 2024a) 28.57 12.20 11.54 25.61 21.23 27.15 18.15 13.39 26.96 18.40 20.87
Gemini-2.5-Flash (Comanici et al. 2025) 68.13 25.61 7.69 61.93 72.07 75.17 55.85 49.11 52.17 39.20 55.77
Chart-Specific Models
ChartReasoner-3B (Ours) 58.24 21.95 22.44 56.24 79.61 86.42 66.14 65.18 54.78 48.80 61.27
ChartReasoner-7B (Ours) 72.53 31.71 32.05 65.43 82.96 84.44 71.55 65.18 66.09 53.60 67.06
Table 2: Comparison of ChartReasoner with representative MLLMs on the porposed CRBench, using Chain-of-Thought (CoT) prompting to elicit the models’ intermediate reasoning steps. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
ChartReasoner
We propose ChartReasoner, a MLLM designed to perform visual reasoning for better visual understanding on non-annotated charts. We formally define this task as follows: given a chart image $I$, and a corresponding textual question $Q$, the goal is to derive a numerical answer $A$ with a reasoning procedure $R$. This process can be represented as a mapping function $\mathcal{F}$:
@@
where $I$ is the chart image, $Q$ is the question in text, $R$ is the step-wise reasoning procedure in text, and $A\in\mathbb{R}$ is the numerical answer. The fundamental challenge lies in interpreting non-annotated charts, which requires the model to reason about geometric structures (e.g., axes, scales, positions) to infer values, rather than simply extracting them via text recognition. While the mapping $\mathcal{F}$ inherently performs this reasoning, producing an explicit reasoning chain $R$ that explains the derivation is an optional, secondary objective.
To address this challenge, we propose a two-stage Reinforcement Finetuning (RFT) strategy. This approach is designed to first instill a robust, human-like reasoning framework and then meticulously refine the model’s numerical precision. As illustrated in Figure 2, the RFT pipeline consists of two sequential stages: (1) Reasoning Activation via SFT, which uses supervised fine-tuning to teach the model the structure of reasoning, followed by (2) Reasoning Generalization via RL, which uses reinforcement learning to improve the accuracy and generalizability.
Stage 1: Reasoning Activation via SFT
The initial stage of our pipeline aims to establish a foundational reasoning paradigm. Instead of having the model directly guess an answer, we teach it to adopt a structured, step-by-step thought process that mirrors human analysis. To achieve this, we first generate a high-quality dataset by distilling detailed Chain-of-Thought (CoT) processes from a MLLM with emerging reasoning abilities. Fine-tuning our base MLLM on these 68,500 synthesized samples teaches the model to emulate a human’s analytical approach, guiding it to identify and utilize critical chart components—such as axes, scales, and legends—before arriving at a final answer.
Formally, we employ SFT on this dataset. Each data instance is a tuple $(x,q,r,a)$, where $x$ is the chart image, $q$ is the question, $r$ is the intermediate reasoning chain, and $a$ is the final answer. The training objective is to maximize the log-likelihood of the model generating the complete sequence $y$ (the concatenation of $r$ and $a$) given the image $x$ and question $q$:
@@
Stage 2: Reasoning Generalization via RL
Building on the reasoning foundation from Stage 1, the second stage focuses on enhancing the model’s precision and reliability for the numerical estimation task. For this, we use a smaller, high-signal dataset of 3,400 samples curated to target the model’s specific weaknesses. These samples are identified by selecting problems where the SFT-tuned model exhibits ”stochastic correctness”—that is, problems it can solve but not consistently. By training on these borderline cases with higher-resolution images, we force the model to refine its visual interpretation skills.
We employ GRPO (Shao et al. 2024), an efficient and scalable reinforcement learning algorithm, to fine-tune the policy $\pi_{\text{SFT}}$. Unlike traditional algorithms like PPO (Schulman et al. 2017), GRPO forgoes a computationally expensive value network and instead calculates relative advantages by comparing rewards within a group of sampled responses. For each input $(x,q)$, we sample a group of $G$ candidate answers $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\beta}$. Each answer $a_{i}$ receives a reward $R(a_{i})$, and these rewards are used to compute a normalized relative advantage $A_{i}$ for each sample:
@@
Reward Function Design
@@
To encourage interpretable and well-structured outputs, we provide a binary format reward, $R_{\text{format}}$. The model receives a reward of 1 if its response strictly adheres to our predefined template—reasoning enclosed in <think></think> tags and the final answer in <answer></answer> tags—and 0 otherwise.
Accuracy Reward.
@@
Based on a human-calibrated tolerance threshold of $\varepsilon=0.02$, we define the reward using a piecewise quadratic function that smoothly decays from 1 to 0:
$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\varepsilon}\right)^{2}&\text{if }d_{\text{rel}}<\varepsilon\\ 0&\text{if }d_{\text{rel}}\geq\varepsilon\end{cases}$ (5)
@@
Figure 3: Qualitative analysis on CRBench. GPT-4o and Qwen2.5-VL-7B exhibits hallucination in the answer and reasoning procedure, respectively. In contrast, our ChartReasoner-7B is able to produce a coherent and correct step-by-step reasoning process, leading to an accurate answer.
Methods CRBench CharXiv (R) CharXiv (D) ChartBench ChartQAPro
ChartGemma-3B (Masry et al. 2025b) 26.74 12.50 21.30 - 6.84
TinyChart-3B (Zhang et al. 2024) 35.83 8.30 16.15 - 13.25
ChartInstruct-7B (Masry et al. 2024) 20.91 8.80 21.40 - 4.88
ChartVLM-7.3B (Xia et al. 2025) 33.63 - - 12.06 -
ChartLlama-13B (Han et al. 2023) 5.01 14.20 19.23 21.30 -
ChartReasoner-7B (Ours) 67.06 43.40 63.90 45.34 36.06
@@
Table 3: Performance of ChartReasoner compared to other chart-specific models on various public chart understanding benchmarks. The results of the compared methods are from the corresponding benchmarks. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
Experiments
Experimental Setups
Training Setting
The implementation was built upon the ModelScope SWIFT framework (Zhao et al. 2025) for all training procedures. We initialize ChartReasoner with Qwen2.5-VL (Bai et al. 2025) and tune it with the dataset described in the previous section. All experiments were carried out on a server with 8 NVIDIA H800 GPUs. The methods evaluated on CRBench follow their default inference settings. All results were obtained from a single forward pass with a fixed random seed of 42. Please refer to the supplementary materials for more details.
Training Paradigms and Baselines
To systematically evaluate our training strategies, we conduct a comparative study in four distinct paradigms.
  • CoT-SFT: Provides a reasoning “cold start” by fine-tuning the model on CoT dataset, with the reasoning data distilled from Qwen2.5-VL-32B (Bai et al. 2025).
  • GRPO-Zero: An ablation study in which our GRPO algorithm is applied directly to the base model to assess the efficacy of reinforcement learning in isolation.
  • RFT: Our proposed final paradigm starts from the CoT-SFT checkpoint and uses GRPO to further refine and generalize the model’s reasoning capabilities.
To further validate the generalization and effectiveness of our training paradigms, we conduct experiments across a diverse set of base models. Specifically, we evaluate our methods on both the Qwen2.5-VL-Instruct (Bai et al. 2025) series, comparing the 3B and 7B variants to assess scalability.
Methods Accuracy
Synthetic Real Overall
ChartReasoner-3B
Zero-Shot 46.41 42.90 45.90
+CoT-SFT 51.02 45.45 50.22
+GRPO-Zero 56.59 56.82 56.62
+RFT 60.92 54.83 60.05
ChartReasoner-7B
Zero-Shot 60.54 45.17 58.34
+CoT-SFT 63.59 51.70 61.88
+GRPO-Zero 66.44 54.83 64.78
+RFT 66.97 60.23 66.00
Table 4: Ablation study showing the individual contribution of each strategy (CoT-SFT, GRPO-Zero, and RFT) to the final performance of the ChartReasoner models. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
Experimental Results
Evaluation on CRBench
The results in Table 1 underscore the significant challenge our CRBench presents to current models. The generally low scores across the board reveal a critical weakness: the widespread inability to reason over non-annotated charts. This difficulty stems from our benchmark’s emphasis on genuine visual reasoning—inferring values from axes and scales—rather than relying on annotation-based shortcuts common in other benchmarks. Furthermore, while most methods obtains comparable results on synthetic and real charts, the overall scores are still unsatisfying. To reveal the failure stems, we employed CoT prompting to compel models to externalize their reasoning process. The results in Table 1 and Table 2 show that most methods do not benefit from CoT prompting. It can be attributed to the lack of visual reasoning ability, as shown in Figure 3.
Evaluation of ChartReasoner
Our comprehensive evaluation confirms that ChartReasoner-7B decisively outperforms not only existing chart-specific models but also top-tier proprietary models like GPT-4o and Gemini-2.5-Flash, as shown in Table 2. This success stems from its ability to perform genuine visual reasoning, a skill that other models lack, as evidenced by their reliance on taking annotation shortcuts. This advanced capability is particularly evident in its handling of complex chart typologies. For instance, on challenging radar charts, where many models falter, ChartReasoner demonstrates a significant performance leap, proving it can interpret complex visual structures.
Ablation Analysis of the Training Paradigm
Our experimental results, summarized in Table 4, strongly validate the effectiveness of the RFT strategy for enhancing visual reasoning. As discussed in the previous section, simply prompting a model with a CoT format often leads to performance degradation, as it forces the model to articulate a reasoning process it does not possess. Our training paradigms were designed to address this fundamental issue. The result clearly indicates that while a naive CoT-SFT offers only a marginal improvement, it is the synergistic combination within the RFT model that yields the most significant performance gains, consistently outperforming all other configurations.
Case Study of Visual Reasoning

### v2

@@ 3.2 Evaluation Protocol
4 ChartVR
We propose ChartVR, a series of MLLMs designed to perform visual reasoning for better visual understanding on non-annotated charts. We formally define this task as follows: given a chart image $I$, and a corresponding textual question $Q$, the goal is to derive a numerical answer $A$ with a reasoning procedure $R$. This process can be represented as a mapping function $\mathcal{F}$:
@@ 4 ChartVR
where $I$ is the chart image, $Q$ is the question in text, $R$ is the step-wise reasoning procedure in text, and $A\in\mathbb{R}$ is the numerical answer. The fundamental challenge lies in interpreting non-annotated charts, which requires the model to reason about geometric structures (e.g., axes, scales, positions) to infer values, rather than simply extracting them via text recognition.
To address this challenge, we propose a novel two-stage Reinforcement Finetuning (RFT) framework. This approach is designed to first instill a robust, human-like reasoning framework and then meticulously refine the model’s numerical e precision. As illustrated in Figure 2, the RFT pipeline consists of two sequential stages: (1) Visual Reasoning Activation, which uses supervised fine-tuning to teach the model the structure of reasoning, followed by (2) Visual Reasoning Generalization, which uses reinforcement learning to improve the accuracy and generalizability.
4.1 Stage 1: Visual Reasoning Activation
The initial stage of our pipeline aims to establish a foundational reasoning paradigm. Instead of having the model directly guess an answer, we teach it to adopt a structured, step-by-step thought process that mirrors human analysis. To achieve this, we fine-tune our base model on a high-quality dataset of 43k samples generated by distilling detailed Chain-of-Thought (CoT) processes from an advanced MLLM (see Appendix B.1 for details). This CoT-SFT process systematically teaches the model to move beyond direct answer prediction and instead adopt a structured analytical approach: first identifying and utilizing critical chart components—such as axes, scales, and legends—and then using them to derive a final answer.
Formally, we employ SFT on this dataset. Each data instance is a tuple $(x,q,r,a)$, where $x$ is the chart image, $q$ is the question, $r$ is the intermediate reasoning chain, and $a$ is the final answer. The training objective is to minimize the negative log-likelihood of the model generating the complete sequence $y$ (the concatenation of $r$ and $a$) given the image $x$ and question $q$:
@@ 4.1 Stage 1: Visual Reasoning Activation
4.2 Stage 2: Visual Reasoning Generalization
Building on the visual reasoning foundation from Stage 1, the second stage focuses on enhancing the model’s precision and reliability for the numerical estimation task. For this, we use a smaller, high-signal dataset of 3.4k samples curated to target the model’s specific weaknesses. These samples are identified by selecting problems where the SFT-tuned model exhibits “stochastic correctness”—that is, problems it can solve but not consistently (see Appendix B.3 for details). By training on these borderline cases with higher-resolution images, we force the model to refine its visual interpretation skills.
We employ GRPO (shao2024deepseekmath), an efficient and scalable reinforcement learning algorithm, to fine-tune the policy model $\pi_{\text{SFT}}$. Unlike traditional algorithms like PPO (schulman2017proximalpolicyoptimizationalgorithms), GRPO forgoes a computationally expensive value network and instead calculates relative advantages by comparing rewards within a group of sampled responses. For each input $(x,q)$, we sample a group of $G$ candidate answers $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\beta}$. Each answer $a_{i}$ receives a reward $R(a_{i})$, and these rewards are used to compute a normalized relative advantage $A_{i}$ for each sample:
@@ 4.2 Stage 2: Visual Reasoning Generalization
4.3 Reward Function Design
@@
To encourage interpretable and well-structured outputs, we provide a binary format reward, $R_{\text{format}}$. The model receives a reward of 1 if its response strictly adheres to our predefined template, where reasoning is enclosed in <think></think> tags and final answer in <answer></answer> tags, and 0 otherwise.
Continuous Accuracy Reward.
@@
Then, we define the reward using a piecewise quadratic function that smoothly decays from 1 to 0:
$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\tau}\right)^{2}&\text{if }d_{\text{rel}}<\tau\\ 0&\text{if }d_{\text{rel}}\geq\tau\end{cases}$ (5)
We empirically set $\tau=0.02$ based on a human-calibrated tolerance threshold. For the specific case where the ground truth $A_{\text{gt}}$ is zero, because it is difficult to quantize the relative deviations, the accuracy reward falls back to exact match, assigning a value of 1 when $A_{\text{gt}}=A_{\text{pred}}$ and 0 otherwise.
@@
Table 2: Comparison of ChartVR with representative MLLMs on the proposed ChartVRBench. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
Methods Synthetic Charts Real Charts Overall
Box Area Radar Scatter Bar Line Combo Bar Line Combo
Human Evaluation 94.51 43.29 88.46 91.24 96.65 97.68 90.92 84.82 84.35 65.60 87.57
Open-source Models
InternVL3-2B (zhu2025internvl3) 25.27 8.54 9.62 25.16 51.68 43.05 34.55 43.75 34.78 32.00 32.98
Qwen2.5-vl-3B (bai2025qwen2) 46.15 14.02 17.95 51.42 72.91 81.13 62.83 66.96 54.78 45.60 56.62
Ovis1.6-llama3.2-3B (lu2024ovis) 12.09 3.66 7.69 14.66 13.13 10.93 11.69 8.04 16.52 12.00 11.66
Gemma-3-4B (team2025gemma) 9.89 3.05 12.82 11.82 9.22 12.25 6.81 8.04 12.17 7.20 9.34
Qwen2.5-vl-7B (bai2025qwen2) 70.33 21.34 19.23 61.93 73.74 85.43 68.41 49.11 56.52 48.00 61.39
InternVL3-8B (zhu2025internvl3) 36.73 12.80 12.18 39.17 43.58 47.68 36.82 38.39 46.09 34.40 36.73
Close-source Models
GPT-4o (hurst2024gpt) 28.57 12.20 11.54 25.61 21.23 27.15 18.15 13.39 26.96 18.40 20.87
Gemini-2.5-Flash (comanici2025gemini) 68.13 25.61 7.69 61.93 72.07 75.17 55.85 49.11 52.17 39.20 55.77
Chart-Specific Models
ChartGemma-3B (masry2024chartgemma) 10.99 10.98 7.05 21.44 42.74 32.78 24.43 37.50 42.61 28.80 26.74
TinyChart-3B (zhang2024tinychart) 13.19 7.93 7.69 25.16 56.15 54.30 36.65 57.14 40.87 34.40 35.83
ChartInstruct-7B (masry2024chartinstruct) 10.99 1.22 4.49 16.63 35.47 16.56 18.85 51.79 45.22 18.40 20.91
ChartVLM-7.3B (xia2024chartx) 9.89 10.37 7.69 10.28 70.39 45.36 32.81 50.00 54.78 35.20 33.63
ChartLlama-13B (han2023chartllama) 10.99 1.83 5.77 5.25 3.35 4.30 4.01 8.04 9.57 7.20 5.01
Bespoke-MiniChart-7B (liyan2025minichart) 72.53 26.83 25.00 66.74 86.87 89.40 75.04 69.64 62.61 45.60 68.16
Chart-R1 (7B) (chen2025chartr1) 83.52 26.22 25.64 66.83 85.47 82.45 68.94 61.61 59.13 44.80 65.72
ChartVR-3B (Ours) 63.74 24.39 19.87 57.33 82.68 87.75 69.46 58.93 57.39 45.60 62.74
ChartVR-7B (Ours) 83.52 37.20 27.56 70.90 88.27 92.05 78.53 69.64 62.61 58.40 72.20
5 Experiments
5.1 Experimental Setups
Implementation Details.
The implementation was built upon the ModelScope SWIFT framework (zhao2025swift). We initialize our ChartVR models using the open-source Qwen2.5-VL series (bai2025qwen2) as a foundation. For inference, all models and benchmarks follow their provided settings where available, with results obtained from a single forward pass using a fixed random seed of 42 to ensure reproducibility. Additional details are available in Appendix C.
Main Evaluation on ChartVRBench.
Our primary evaluation is conducted on the proposed ChartVRBench to assess genuine visual reasoning capabilities and establish the superiority of our ChartVR model. On this benchmark, we compare our model against a comprehensive suite of baselines organized into three categories: open-source MLLMs, powerful close-source MLLMs and prominent chart-specific models.
Generalization Study on Public Benchmarks.
To evaluate the transferability of the skills learned via our RFT framework, we conduct a generalization study. Specifically, we benchmark our ChartVR model against the representative chart-specific models—ChartGemma (team2025gemma), TinyChart (zhang2024tinychart), ChartInstruct (masry2024chartinstruct), ChartVLM (xia2024chartx), and ChartLlama (han2023chartllama)—on a diverse set of public benchmarks. This suite includes the real-world datasets CharXiv (wang2024charxiv) and ChartQAPro (masry2025chartqapro), as well as the synthetic benchmark ChartBench (xu2023chartbench). This allows us to verify that our training methodology imparts a foundational reasoning ability that generalizes effectively to a variety of chart understanding tasks.
5.2 Experimental Results
@@
Table 3: Performance of ChartVR compared to other chart-specific models on various public chart understanding benchmarks. All results are reproduced by the authors. Qwen2.5-VL baselines are listed below their corresponding ChartVR models with an arrow indicator. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
Models ChartVRBench CharXiv (R) ChartBench ChartQAPro
ChartGemma-3B (masry2024chartgemma) 26.74 12.50 - 6.84
TinyChart-3B (zhang2024tinychart) 35.83 8.30 - 13.25
ChartInstruct-7B (masry2024chartinstruct) 20.91 8.80 - 4.88
ChartVLM-7.3B (xia2024chartx) 33.63 - 12.06 -
ChartLlama-13B (han2023chartllama) 5.01 14.20 21.30 -
Bespoke-MiniChart-7B (liyan2025minichart) 68.16 41.40 44.19 34.74
Chart-R1 (7B) (chen2025chartr1) 65.80 50.00 15.71 31.81
ChartVR-3B (Ours) 62.74 33.40 26.35 28.03
   $\hookrightarrow$ Qwen2.5-VL-3B (bai2025qwen2) 56.62 30.60 21.30 24.51
ChartVR-7B (Ours) 72.20 43.40 45.34 41.79
   $\hookrightarrow$ Qwen2.5-VL-7B (bai2025qwen2) 61.39 39.50 35.35 37.10
Performance on ChartVRBench.
The results, presented in Table 2, underscore the significant challenge that ChartVRBench poses to a wide range of MLLMs. The generally low scores across all categories—including powerful proprietary models like GPT-4o (20.87%) and Gemini-2.5-Flash (55.77%)—reveal a critical and widespread weakness in genuine visual reasoning. This difficulty stems from our benchmark’s design, which forces models to infer values from graphical geometry (e.g., axes and scales) rather than relying on OCR-based shortcuts common in other benchmarks. The particularly low score of GPT-4o has been double-checked and manually validated, which precisely indicates that many MLLMs lack the genuine visual reasoning ability, as we argue.
Our proposed model, ChartVR, demonstrates a clear superiority in this demanding task. ChartVR-7B achieves an overall score of 72.20%, outperforming all other models, including the best open-source baseline, Qwen2.5-vl-7B (61.39%), and the strongest proprietary model, Gemini-2.5-Flash. Notably, even our smaller ChartVR-3B model (62.74%) surpasses most other models, highlighting the effectiveness of our training methodology. The performance is particularly strong on complex chart types like Real Combo charts, where ChartVR-7B (58.40%) dramatically outperforms the other models.
Generalization on Public Benchmarks.
As detailed in Table 3, ChartVR-7B exhibits exceptional generalization, achieving competitive results across the board. Specifically, on the reasoning-focused portion of CharXiv, our model achieves an improvement of 3.9% over the base model Qwen2.5-VL-7B. This success stems from our model’s core visual reasoning capability, which contrasts with other chart-specific models that rely on SFT and often fail to develop a generalizable reasoning capability.
Table 4: Ablation study of training strategies for ChartVR-3B and ChartVR-7B models. The best and second-best scores are highlighted.
Training Paradigm ChartVR-3B ChartVR-7B
Synthetic Real Overall Synthetic Real Overall
Zero-Shot 56.87 54.78 56.52 67.66 51.16 61.39
+CoT-SFT 36.55 34.86 40.77 55.45 45.74 54.06
+GRPO 56.59 56.82 56.62 66.44 54.83 64.78
+RFT 64.26 53.69 62.74 73.68 63.35 72.20
Table 5: Ablation study of reward function components for Qwen2.5-VL-7B.
Reward Component Score
Format 61.84
Cont. Acc 67.88
Acc + Format 70.28
Cont. Acc + Format 72.20
5.3 Ablation Study
Effectiveness of the RFT Training Paradigm.
To systematically validate our training strategies, we conducted an comparative study comparing three paradigms: CoT-SFT, GRPO applied directly to the base model, and our full RFT framework. The results are summerized in Table 4. For ChartVR-7B, the base model scores 61.39%. Using CoT data as a ’Visual Reasoning Activation’ merely compels the model to adopt a visual reasoning pattern without imparting the underlying ability. Consequently, this mismatch leads to a performance degradation rather than an improvement. In contrast, the full RFT framework, which synergistically combines SFT with RL, achieves the most significant performance gain, reaching 72.20%. This demonstrates that our complete VR-RFT framework genuinely enhances the model’s visual reasoning ability, leading to its consistent and substantial outperformance over all other configurations.
Impact of the Reward Function Design.
We conducted an ablation study to isolate the contribution of each component in our reward function, with results presented in Table 5. The findings highlight a strong synergistic effect between enforcing a correct output structure and rewarding numerical precision. A model trained with only the Format reward achieves a score of 61.84, whereas combining it with our proposed Continuous Accuracy (Cont. Acc) reward boosts the score significantly to 72.20, indicating both components are crucial. Furthermore, the study validates the superiority of our continuous reward design over a standard binary alternative. A model using a simple binary accuracy reward (Acc + Format), which provides a sparse correct/incorrect signal, is clearly outperformed by our model using the continuous reward (Cont. Acc + Format). This demonstrates the effectiveness of our quadratic reward function, which provides a dense and informative learning gradient. By rewarding ”nearly correct” answers, it encourages the fine-grained improvements in visual estimation necessary for achieving higher precision.
5.4 Case Study
Figure 3: Qualitative analysis on ChartVRBench. GPT-4o and Qwen2.5-VL-7B exhibits hallucination in the answer and reasoning procedure, respectively. In contrast, our ChartVR-7B is able to produce a coherent and correct step-by-step reasoning process, leading to an accurate answer.